M 362K 60165 Fall ’06 T Th 11-12:30 RLM 6.116

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30 or by appt. Web Page (homework assignments are listed): www.ma.utexas.edu/˜ friedman

Text and Syllabus: Introductory Probability and Statistical Applications , 2nd ed. Meyer, Addison Wesley

(This will be published by the University Co-op; the original text is out of print.)

Syllabus: Parts of Chapters I-XIV.( Exactly which sections will be covered, depends on how quickly we progress, etc.)

Prerequisites: M408C,D or M808A,B including infinite series, partial derivatives, double integrals.

Exams: We will have exams on 10/3 (class 10; Tues.), 11/7 (class 20; Tues.), 12/5 (Class 27; Tues.) No “Final exam” (subject to Dept. approval.).

* Exam grades will be scaled (if the unscaled average is ≤ 75) to some value in the 75 to 80 range. ** Makeups will be given very reluctantly.

Homework: Homework problems will be assigned each Tuesday in class and will be due the next Tuesday. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework together will be counted as a grade equivalent to one exam grade.

Final Grades: In computing a final grade, I will count the three tests (the Final is one of them) and one total homework grade equally. (So there will be four grades to average.)